OpenStudy (kaiiskool):

Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2 will fan and medal.

OpenStudy (daniel.ohearn1):

You have options for your answer?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (daniel.ohearn1):

If y will be in the form y=mx+b you can get this answer by multiplying your x value by 2

OpenStudy (kaiiskool):

y + 9 = 2(x – 5) y – 9 = 2(x – 5) y – 5 = 2(x + 9) y – 9 = 2(x + 5)

OpenStudy (daniel.ohearn1):

m=2

pooja195 (pooja195):

@daniel.ohearn1 it says point slope form also please let the asker do the work!

OpenStudy (kaiiskool):

alright guys thanks ill fan both of you for all the help

pooja195 (pooja195):

xD ok well we still need to finish this :P

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (kaiiskool):

yeah i just needed the work youve guys have helped enough

pooja195 (pooja195):

oh ok yw